INTRAMOLECULAR DYNAMICS IN X_{4} RHOMBIC CLUSTERS

\underline{ROBERTO LINGUERRI}, ISABELLE NAVIZET, and PAVEL ROSMUS, Laboratoire de Chimie Théorique, Université de Marne la Vallée, F 77454 - Champs sur Marne, France; STUART CARTER, Department of Chemistry, University of Reading, Reading RG6 2AD, U.K.; JOHN P. MAIER, Department of Chemistry, University of Basel, Klingelbergstr. 80 CH-4056 Basel, Switzerland.

Two independent approaches are applied to study the intra-molecular dynamics in B_{4} electronic ground state. First, a double minimum six-dimensional potential energy surface (PES) is determined in symmetry coordinates for the most stable rhombic (D_{2h}) B_{4} isomer in its ^{1}A_{g} electronic ground state by fitting to energies calculated by the CCSD(T) approach. The PES exhibits a barrier to the D_{4h} square structure of 255 cm^{−1}. The vibrational levels (J = 0) are calculated variationally using the MULTIMODE^{a} code, which involves the Watson kinetic energy operator expressed in normal coordinates. The pattern of about 65 vibrational levels up to 1600 cm^{−1} for all stable isotopomers is analyzed. Analogous to the inversion in ammonia-like molecules, the rhombus rearrangements lead to splittings of the vibrational levels. In B_{4} it is the b_{1g} (D_{4h}) mode which distorts the square molecule to its planar rhombic form. It is shown that the rhombic ground states of B_{4}^{+} and B_{4}^{−} consist of two vibronically coupled electronic states. The PES for B_{4} is used also in time-dependent molecular dynamics calculations using the multi-configurational time-dependent Hartree^{b} (MCTDH) method. The 6-D wavepackets are analyzed and the results are compared with time-independent results.

aS. Carter, S. Culik and J. M. Bowman J. Chem. Phys. \underline{107}, 10458 (1997).

bH.-D. Meyer, U. Manthe, and L. S. Cederbaum. Chem. Phys. Lett. \underline{165}, 73 (1990).